An object is thrown straight upward from the ground with an initial speed of 30 feet per second. Its height (h) after t seconds is given by the equation h(t) =
-5t2 + 30t. Find the domain for h(t).
A. (-∞, 0)
B. (6, ∞)
C. (0, 6)
D. All real numbers

Answer by stanbon(75887)

An object is thrown straight upward from the ground with an initial speed of 30 feet per second. Its height (h) after t seconds is given by the equation h(t) =
-5t2 + 30t. Find the domain for h(t).
A. (-∞, 0)
B. (6, ∞)
C. (0, 6)
D. All real numbers
------
Solve: -5t^2+30t = 0
-5t(t-6) = 0
t = 0 or t = 6
-------------------------
As a function the Domain is "All Real Numbers".
----
As an application to the physical problem the domain is (0,6)
===============
Cheers,
Stan H.